The equation of the straight line passing through the point (4,3) and making intercepts on the coordinate axes whose sum is (-1), is-
A) 3x+2y=6 and x+2y=2
B x/2-y/3=1 and x/(-2)+y/1=1
C) x/2+y/3= -1 and x/(-2)+y/1=-1
D x/2-y/3=-1 and x/(-2)+y/1=-1

Grade:11

1 Answers

Arun
4 years ago

As we learnt in

Intercept form of a straight line -

\frac{x}{a}+\frac{y}{b}=1

 

- wherein

a and b are the x-intercept and y -intercept respectively.

 

 Let the equation of straight line be 

\frac{x}{a}+\frac{y}{b}=1

i.e.  \frac{4}{a}+\frac{3}{b}=1                                                ..........(1)

also\ a+b=-1                                         ...........(2)

4b+3a=ab

4(-1-a)+3a=a(-1-a)

-4 - a = - a - a2

\Rightarrow a^{2}=4\, \, \, \, \Rightarrow a=\pm2

and a+b=-1

if a=2, b=-3

a=-2, b=1